There is a uniform electrostatic field in a region. The potential at various points on a small sphere centred at P, in the region, is found to vary between the limits 589.0V to 589.8V. What is the potential at a point on the sphere whose radius vector makes an angle of 600 with the direction of the field ?
(1)589.5V(2)589.2V(3)589.4V(4)589.6V
1
Expert's answer
2019-08-01T09:54:08-0400

Potential gradient is equal to


"v=589.8 V \u2013 589.0 V = 0.8 V (1)"

Potential gradient is given by formula:


"v=Ed \cos 60° (2)"

Using (1) and (2) we got:


"v=0.8 V\u00d70.5 =0.4 V (3)"

The potential at a point on the sphere is equal to


"v=589.0 + 0.4 V= 589.4 V"

Answer:

(3)589.4V